Public Policy 246 The Community Reinvestment Act

Dustin Mitchell

March 26, 2001

1 Introduction

Racial discrimination has been a pervasive force in housing since the second world war. Based on a study by University of Chicago researcher Homer Hoyt correlating race to changes in property values, the Federal Housing Authority itself explicitly encouraged segregation with its lending policies: “If a neighborhood is to retain stability, it is necessary that properties shall continue to be occupied by the same social and racial classes. A change in social or racial occupancy generally contributes to instability and a decline in values.”^{1} This attitude was reflected by private lenders, resulting in a situation where lenders refuse to extend credit to low-income and/or minority areas, often inner cities. This type of lending discrimination is called redlining, and is a powerful force in furthering disinvestment in the inner cities.^{2}

Although they are not served by creditors, low- and middle-income individuals do make use of bank services, and in particular demand-deposit accounts. In many cases banks accept deposits from disadvantaged areas, but neither lend, invest, nor provide adequate bank services (branch offices, ATMs, account features) in those areas. This practice has served as a conduit to siphon funds from inner cities and invest them elsewhere.

1U.S. Federal Housing Administration 1938, par. 937; cited in Revitalizing urban neighborhoods, ed. W. Dennis Keating, Norman Krumholz, and Philip Star (Lawrence, Kansas:

University Press of Kansas, 1996), 223.

2Revitalizing urban neighborhoods, 22.

Beginning in the late 1960’s, Congress took action to change this situation. The Federal Fair Housing Act (1968) and the Equal Credit Opportunity Act (1968) outlawed explicit discrimination. The Home Mortgage Disclosure Act (1975) required lenders to make publicly available the geographic distribution of their loans. Finally, the Community Reinvestment Act (1977) mandated that banks properly serve their deposit areas.

The goal of the CRA is to encourage banks to meet the credit needs of the communities from which they accept deposits.^{3} In particular, institutions must lend, invest, and provide services in areas from which they accept deposits.^{4} However, the authors of the CRA were careful to point out that the new policies need not and should not conflict with safe and sound banking operations.

2 Implementation

The Community Reinvestment Act is implemented as a part of the preexisting examination which is done by the federal government of all lending institutions. In addition to its more usual ratings, a bank’s efforts to meet the credit needs of its communities is scored on a four-point scale.^{5}

Specifically, the institutions required to file CRA data are large state banks, national banks, and savings associations. Evaluation and enforcement of the CRA is done jointly by four federal agencies: The Office of Thrift Supervision (OTS), the Office of the Comptroller of the Currency (OCC), the Board of Governors of the Federal Reserve System (FRB), and the Federal Deposit Insurance Corporation (FDIC).^{6} The criteria used in these evaluations include the quality of the bank’s efforts to assess the needs of its communities and to advertise its services to potential customers, as well as the type and geographic distribution of offered credit. Institutions are also evaluated on their involvement in community development practices, particularly with regards housing, small business, and (in rural areas) farming. Beyond these guidelines, there are no explicit standards set forth in the CRA, and considerable freedom is left to banks to find ways to meet their responsibilities.^{7}

3FFIEC Community Reinvestment Act: Background & Purpose, Federal Financial Institutions Examination Council, 24 August 1999, <http://www.ffiec.gov/cra /backpur.htm>.

4NAMIC - Community Reinvestment Act, National Association of Mutual Insurance Companies, <http://www.namic.org/cra/3.htm>.

5Inner City Press’ Community Reinvestment Act Guide & News, Inner City Press, <http://www.innercitypress.org/cra.html>.

A bank’s performance rating is taken into account when it petitions for federal approval for deposit insurance, a new branch, a merger, or an acquisition. These petitions can also be protested by third parties. In many cases it is through these protests that lending institutions are forced to make a substantive commitment to meet the needs of their communities. Often institutions will, with the encouragement of the regulators, make voluntary agreements with protesters, resulting in the withdrawal of the protest.^{8}

3 Case Studies

We will now very briefly examine several cases in which the CRA was applied, with varying levels of success.

6FFIEC Community Reinvestment Act: Background & Purpose

7Community Reinvestment Act 1977-1987: Tenth Anniversary Forum, Columbia, South Carolina, 28 October 1987, ed. James F. Tucker (Richmond: The Federal Reserve Bank of Richmond, 1988), 7.

8Revitalizing urban neighborhoods, 226.

3.1 Chicago

In 1984, the Community Reinvestment Alliance, a collective of Chicago community groups, protested petitions to the FRB by three banks — Northern Trust Co., American National Corp., and Bank of Montreal. The alliance and the banks met and negotiated mutually agreeable terms, and the protests were dropped. The agreement included extensions of credit to residential, commercial, and industrial areas with median incomes below 75% of the Chicagoland median income. Although loan prices are often below the usual market prices, they are not below cost, and thus not unprofitable for banks.

For its part, the alliance took on the role of loan packager: certain organizations act as a front-end for the bank, advertising the availability of credit, doing preliminary screening of applicants, and assembling the necessary documentation and paperwork to ease the processing required from the bank. In order to continue their partnership with the community groups, each of the subject banks has established a review board to oversee the implementation of the agreements, and to fine-tune the terms as the economic environment changes.

The community groups are very pleased with the progress of their partnership. In addition to the rise in low- and middle-income loans made by the partner banks, community groups were pleased to see multi-family mortgages to low- and middle-income areas nearly double. The banks are also happy with the situation. Not only do they have the assistance of community groups in the administrative work of meeting their CRA requirements, but the loans they have made have performed very well. For example, in 2.5 years, Northern Trust lent $11 million with no defaults nor even a missed payment.^{9}

9“From Protests to Partnerships,” ABA Banking Journal, September 1987, 40-44.

3.2 ACORN in St. Louis

Another success story comes from Boatman’s Bank in St. Louis. This bank’s holding company was making almost no loans to the needy areas of St. Louis, preferring instead to focus on the more affluent suburbs and outlying towns. When another company applied to acquire the holding company, ACORN (the Association of Community Organizations for Reform Now) protested. In the negotiations that ensued, Boatman’s agreed to invest $50 million in St. Louis and to establish a review committee to oversee the investment.^{10}

Boatman’s officials are enthusiastic about the agreement. Says Ethan A.H. Shepley, Jr., “We’re being held up as the heroes, and that’s super. That’s the kind of publicity we love. And we’re not making any loans we feel are bad.”^{11}

Here again is evidence that the CRA encourages banks to make safe, sound commitments of credit to areas that are excluded by established lending policies.

3.3 ACORN and Hibernia National Bank

In this well-publicized case, ACORN protested the acquisition of a Louisiana bank by Hibernia National Bank in New Orleans. Less than a year before the petition for acquisition, the OCC had given the bank a “2” for CRA compliance – satisfactory. However, the FRB performed an examination in response to ACORN’s protest, and found the bank’s performance woefully inadequate. Indeed, Hibernia had not even kept the proper records to support an evaluation.

10Mildred Brown, written testimony to United States Senate, Committee on Banking, Housing, and Urban Affairs, The Implementation and Enforcement of the Community Reinvestment Act of 1977 (Washington: GPO, 1988), 23-24.

11Brown, 24.

After an eight-month delay, the FRB approved the acquisition, conditioned on Hibernia developing and submitting a plan to improve its performance. The FRB also stated that no new petitions by Hibernia would be approved until certain substantive improvements had been made. However, the FRB would not, as a matter of policy, release any of the documentation of Hibernia’s improvements to ACORN. Despite this, ACORN soon learned that Hibernia failed to submit any of the necessary documents, and that several more petitions from Hibernia were approved by the FRB shortly thereafter.

It is clear that this case, unlike the previous two, represents a catastrophic failure in the enforcement of the CRA. Although Hibernia’s first application was delayed, at significant expense, the bank did not make any substantive changes in its policy toward its low- and middle-income deposit areas.^{12}

4 Benefits

It is clear from the examples that the CRA can and does work well in many cases. However, government involvement in these successes is often minimal. In both the St. Louis and Chicago examples, it was community groups like ACORN and, in Chicago, NTIC (the National Training and Information Center) that did the work of convincing the banks to make commitments to low- and middle-income neighborhoods. This gives focus to the community groups and often (as was the case in Chicago) brings several groups together to put pressure on one bank.

These new, focused organizations provide services that banks are not easily able to provide — a bank generally does not have the human resources to reach out to a low-income community and evaluate its needs, nor can it effectively and cheaply screen and educate potential customers. These tasks are ideally suited to small non-profits which are often community-based and thus have the trust of the community.

12Brown, 17-21.

Overall numbers for the performance of CRA in the last 22 years are not readily available, but many of the local figures are positive. For instance, Chicago home mortgages rose by 50% from 1990-94. In low- and middleincome areas, the increase was 85%. While there were certainly many factors contributing to this change, the work of community groups and banks to build lending in needy areas has certainly been a strong factor.^{13}

5 Issues

Although the CRA has been successful in many cases, it is not without its drawbacks and issues. The problems fall into two major categories. First, lending institutions and community groups alike have many complaints about the policies and practices of government organizations: in many cases the actions of these organizations hinder the implementation of the CRA as it was intended. Second, it is not clear that, if fully implemented, the CRA would indeed be beneficial to low- and middle-income individuals.

5.1 Government Issues

Community groups’ primary complaint about CRA enforcement is that the periodic evaluations are woefully inadequate. In 1984, the OCC only gave one rating worse than “satisfactory” to any of the 254 large banks they evaluated. In light of the findings of the FRB in the Hibernia case, it is very unlikely that all of these ratings are accurate.^{14}

13Rochelle L. Stanfield, “Home, Sweet Home” The National Journal 28, no. 37 (14 September 1996): 1947+.

14Brown, 21.

Mildred Brown, president of ACORN in 1987, conjectures that OCC’s failure to adequately examine banks community lending record is due to lack of concern on the part of the examiners. CRA examinations are only a small part of the overall investigation of an institution, and do not include such crucial components as an investigation of community credit needs.^{15} The OCC may feel that protests against mergers and acquisitions will adequately identify problematic institutions, and that their own evaluations are virtually irrelevant to the process. This attitude, however, is directly contrary to the letter and the spirit of the Act.

In lodging a protest, community groups are often slighted. The FRB is responsible for investigating all protests, encouraging negotiations between the community and the lending institution, and in concluding and officiating any agreements made in the course of those negotiations. Protests must be submitted during a comment period, which is usually 30 days.^{16} One common problem expressed by community groups is that 30 days is far too short a time to investigate a lending institution and produce a substantive protest. The FRB is often willing to extend the comment period, but in recent years deregulation has reduced this flexibility, sometimes leaving community groups unable to oppose a petition. This change, which hinders enforcement of the CRA by community groups, combined with the OCC’s attitude toward periodic evaluations, seriously diminishes the effectiveness of the CRA.

Full implementation of CRA policy in banks is also hindered by secondary mortgage markets and Private Mortgage Insurers. These are not directly subject to CRA regulation, and the PMIs in particular still practice what amounts to redlining. Both secondary markets and PMIs will not buy loans for properties on blocks with more than 15% abandonment. Further, many institutions require a fixed dollar amount as down-payment, so loans for dwellings affordable to low-income families carry an enormous down-payment when considered as a fraction of the purchase cost.^{17}

15Brown, 21-22.

16Community Reinvestment Act Forum, 8-9.

5.2 Service Delivery Problems

Once banks have negotiated and implemented CRA agreements, there are still obstacles to be overcome. The first of these is that bank administrators often think of their their CRA obligations as writeoffs or grants. As was seen in the Chicago case, however, loans to low- and middle-income neighborhoods are not any riskier than more commonplace loans, nor are they unprofitable. Many banks include grants in their CRA data, but CRA is intended to foster the extension of credit into needy neighborhoods.^{18}

Another significant obstacle to the CRA’s effectiveness is that aggressive marketing of credit to low-income areas often encourages customers to overcommit themselves. For instance, a customer might, in consultation with a lender or loan packager, calculate that she can afford mortgage payments for a new house, but only just barely. After the loan is made and the customer has taken up residence in the house, spontaneous costs arise — the basement floods or the roof requires repair — and the customer is unable to pay these additional costs. Soon, the bank has foreclosed on the property and the customer is left in a worse position than when she started.^{19}

17Brown, 25.

18“From Protests to Partnerships”, 44.

19Elizabeth Berke, City of Cleveland Heights Housing Department, personal conversation, November 1999.

6 Potential Improvements

Although its fundamental intent seems to be effective, there is much discussion of possible modifications to the CRA. Many of these improvements are clear from the discussion above. Clearly there is a need for more explicit guidelines for evaluation of lending institutions. Faced with honest regulatory agencies willing to give them bad marks for bad performance, banks will readily begin the process of community reinvestment usually by allying with a community organization to evaluate the needs of the community. After 22 years, the banking industry has accepted the permanence of the CRA and there is sufficient precedent that a bank can implement an appropriate community lending program with confidence that it will not be too costly.

The information and options available to community organizations could be improved. Specifically, banks and regulators should make more information publicly available, so that community groups can more carefully track banks’ performance on their commitments. Such increased disclosure would have allowed ACORN to force Hibernia to make good on its agreements. In the years since the original bill was passed, amendments have been made to increase the disclosure of CRA information. CRA figures for mortgage lending by individual institutions are now easily available on the Web.

Once a protest has been filed and the FRB has rendered its decision, there is no means for either the bank or any interested community group to appeal the decision. While this has the advantage of preventing institutional petitions from being held up indefinitely, it does exempt the FRB from any direct check on its powerful decisions.

It was mentioned in the discussion of secondary markets and FMIs that these institutions are not within the jurisdiction of the CRA. While such institutions do not accept deposits and thus have no ties to a particular geographic area, there is a clear need for greater regulation in this area. If FNMA and FHLMC are held responsible for the social and racial diversity of the loans they buy, then full implementation of the CRA by banks will no longer require those banks to hold CRA loans in their portfolios.

7 Conclusion

The CRA has been fairly effective in increasing bank lending, investment, and services in low- and middle-income areas. However, much remains to be done. There are very clear improvements which could be made to the act, making it easier to enforce and more appropriately suited to the needs of those using it. The number of protests and active community organizations working with banks is proof that there is still a great amount of work to be done.

The CRA has hardly impacted the larger problem of disinvestment in the inner city. Loans made to inner cities, while growing in number and dollar amount, still do not fully address the financial needs of the communities. The CRA depends on finding the individuals in disinvested communities who are not bad credit risks and who have the personal resources to take out a loan and make good use of it. It is unfortunate, but given the social situation in many American inner cities, such individuals are in a minority. The CRA needs the help of other policies which address the other environmental and social effects of the inner city if it is to be truly effective in revitalizing communities.

8 Works Consulted

Berke, Elizabeth, City of Cleveland Heights Housing Department. Personal conversation. November 1999.

Brown, Mildred. Written testimony to United States Senate, Committee on Banking, Housing, and Urban Affairs. The implementation and enforcement of the Community Reinvestment Act of 1977. Washington: GPO, 1988. Pp. 15-30.

Community Reinvestment Act 1977-1987: Tenth anniversary forum, Columbia, South Carolina, 28 October 1987. Edited by James F. Tucker. Richmond: The Federal Reserve Bank of Richmond, 1988.

FFIEC Community Reinvestment Act: Background & purpose. Federal Financial Institutions Examination Council. 24 August 1999. <http://www.ffiec.gov/cra/backpur.htm>.

“From protests to partnerships.” ABA Banking Journal (September 1987): 40-44.

Inner City Press’ Community Reinvestment Act guide & news. Inner City Press. <http://www.innercitypress.org/cra.html>.

NAMIC - Community Reinvestment Act. National Association of Mutual Insurance Companies. <http://www.namic.org/cra/3.htm>.

Revitalizing urban neighborhoods. Edited by W. Dennis Keating, Norman Krumholz, and Philip Star. Lawrence, Kansas: University Press of Kansas, 1996.

Stanfield, Rochelle L. “Home, sweet home.” The National Journal 28, no. 37 (14 September 1996): 1947+.